ADNAN MUHAMMED

Grade 12,

Calculate the value of gravitational potential energy when a body is moved from the surface of earth to a point ‘h’ above the surface. (I want completely solved answer) plz answer …........ it’s very urgent

Calculate the value of gravitational potential energy when a body is moved from the surface of earth to a point ‘h’ above the surface.
(I want completely solved answer)
plz answer …........ it’s very urgent
 

Grade:12

1 Answers

Aman gohel
6 years ago
=>Acceleration=g(9.8m/s^2)=>mass=m=>Therefore, weight of body or force acted upon body=mg=>Displacement of body=h=>Work done on body=potential energy gained by body=F×S=>W=mg×h=mgh[Ans]